Write in Standard form. Find a: 3x2^ = 2x + 1
a) -3
b) 2
c) 1
d) 3
e) None of these

Answer by stanbon(75887)

3x^2 = 2x + 1
3x^2-2x-1=0
(3x+1)(x-1)=0
x=-1/3 or x=1
Cheers,
Stan H.